List and graph the 6th roots of unity. What are the generators of this group? What are the primitive 6th roots of unity?
Answer #1

Sol from the given data 2,52 Z- 22 23:- from here we know 25- 20= z = (1 tiss/ zz=fltis/ 23 -1 zu = (-1-153) zs= (1-1)/ Hence